# “Quantum Geometric Nesting” and Solvable Model Flat-Band Systems

### v2

@@ Abstract
The properties of a correlated electronic system are largely determined by its electronic structure. One such familiar manifestation is the role of Fermi surface nesting conditions Lomer (1962); Fedders and Martin (1966): if one portion of a Fermi surface overlaps with another under a shift by a wavevector $\bm{Q}$ (i.e. $\bm{k}\to\bm{k}+\bm{Q}$), the Fermi surface is said to be “perfectly nested” and consequently the system is prone to density wave (DW) order at $\bm{Q}$ with particle-hole (p-h) fermion bilinear order. If the overlap occurs under a momentum-space inversion about a certain wavevector $\bm{Q}/2$ (i.e. $\bm{k}\to\bm{Q}-\bm{k}$), the favored order is superconductivity (SC) with particle-particle (p-p) fermion bilinear order and a Cooper pair center of mass momentum $\bm{Q}$. (Such perfect nesting always exists at $\bm{Q}=0$ in time-reversal symmetric systems, where it leads to the familiar Cooper instability.) The thinking behind Fermi surface nesting is based on weak coupling mean-field analysis, in which ordering tendencies are diagnosed by divergent bare susceptibilities in corresponding channels. These concepts are indeed convenient and powerful (see e.g. Refs. Whangbo et al. (1991); Chuang et al. (2001); Kivelson et al. (2003); Shen et al. (2005); Lee et al. (2006); Johannes and Mazin (2008); Terashima et al. (2009); Ran et al. (2009); Nandkishore et al. (2012); Feng et al. (2013); Lee (2014); Jin et al. (2022) for typical applications). However, their validity is limited to weakly coupled systems in which the interaction strength is small compared to the bandwidth. In general, all forms of non-interacting susceptibilities grow as decreasing $T$ with the same asymptotic behavior $\chi\sim 1/T$ for $T$ greater than the bandwidth, which thus indicates comparable ordering tendencies for all possible order parameters. Significantly, this concept, defined solely by the electronic dispersion, overlooks another important piece of information in the electronic structure - the quantum geometry Provost and Vallee (1980); Resta (2011) and topology Bradlyn et al. (2017); Po et al. (2017); Kruthoff et al. (2017) encoded in the single-particle wavefunctions. An entirely different concept of “nesting” is thus needed for flat-band systems, whose bandwidths are small compared to the interaction strengths but which often have non-trivial quantum geometry and rich ordering phenomena Peotta and Törmä (2015); Tovmasyan et al. (2016); Kang and Vafek (2019); Rhim and Yang (2019); Hofmann et al. (2020); Bultinck et al. (2020); Kwan et al. (2021a); Khalaf et al. (2021); Lian et al. (2021); Bernevig et al. (2021); Ledwith et al. (2021); Kwan et al. (2021b); Vafek and Kang (2021); Peri et al. (2021); Liu et al. (2021); Rossi (2021); Angeli and MacDonald (2021); Wang et al. (2022); Hu et al. (2022); Huhtinen et al. (2022); Törmä et al. (2022); Chan et al. (2022a); Herzog-Arbeitman et al. (2022a); Mera and Mitscherling (2022); Pan et al. (2022); Zhang et al. (2022); Herzog-Arbeitman et al. (2022b); Lin et al. (2022); Herzog-Arbeitman et al. (2022, 2022); Peotta (2022); Khalaf et al. (2022); Chan et al. (2022b); Kang and Vafek (2019); Hofmann et al. (2023); Wan et al. (2023); Swaminathan et al. (2023); Hu et al. (2023); Löthman and Black-Schaffer (2017); Chen and Law (2023); Iskin (2023a); Peotta et al. (2023); Iskin (2023b); Tam and Peotta (2023); Zhang et al. (2023); Pan et al. (2023); Crépel et al. (2023); Mao et al. (2023); Kwan et al. (2023); Sahay et al. (2023); Jiang and Barlas (2023); Scheer and Lian (2023a, b); Devakul et al. (2023); Herzog-Arbeitman et al. (2023); Jia et al. (2023); Wu et al. (2023); Chen and Huang (2023); Salerno et al. (2023); Huang et al. (2022); Yu et al. (2023); Kashihara et al. (2023); Soldini et al. (2023); Jiang et al. (2023); Yu et al. (2023); Peralta Gavensky et al. (2023); Setty et al. (2023); Verma et al. (2023); Setty et al. (2023a).
@@
Indeed, except for symmetry-enforced cases, it is unlikely for a physical system to have exact QGN - in common with the case of Fermi surface nesting. However, the use of QGN can be much more versatile: for systems with almost perfect QGN (quantification of the closeness to the ideal nesting will be defined below as ‘nestability’), one may engineer an idealized electronic structure with perfect QGN and then construct ideal interacting Hamiltonians that are close to the physical one. The solvability of these ideal models thus makes them excellent starting points for including perturbative corrections to reach the physical Hamiltonian. This type of intrinsically strong coupling viewpoint (in the sense that the dominant part of the interactions has been accounted for in the solvable models) has proven fruitful in recent studies of twisted bilayer graphene and other flat-band materials Tovmasyan et al. (2016); Kwan et al. (2021b); Vafek and Kang (2021); Kang and Vafek (2019); Khalaf et al. (2021); Bultinck et al. (2020); Lian et al. (2021); Bernevig et al. (2021); Ledwith et al. (2021); Pan et al. (2022); Lin et al. (2022); Herzog-Arbeitman et al. (2022); Zhang et al. (2022); Herzog-Arbeitman et al. (2022b); Crépel et al. (2023); Zhang et al. (2023); Pan et al. (2023); Mao et al. (2023); Kwan et al. (2023). It is noteworthy that perfect QGN is an underlying feature of most of the known examples of solvable models for flat-band systems featuring fermion bi-linear order. We hence expect that QGN can greatly facilitate studies within this paradigm, in a way similar to the roles played by ‘ideal band conditions’ Roy (2014); Claassen et al. (2015); Mera and Ozawa (2021); Ledwith et al. (2022); Wang et al. (2021); Wang and Liu (2022); Wang et al. (2023); Dong et al. (2023) and “vortexability” Ledwith et al. (2023) in studies of fractional Chern insulator states.
@@ I.2 General Definitions
Depending on the channel, momentum $\bm{Q}$, and nesting matrix $N^{\bm{Q}}$, Eq. 8 or 9 represent fermion bilinear orders with different natures. In the p-p channel, if $\bm{Q}=\bm{0}$, the order is a uniform SC order; while for $\bm{Q}\neq\bm{0}$ it is pair-density-wave (PDW, not to be confused with the other DWs in the p-h channel) order of the single-$\bm{Q}$, Fulde–Ferrell type Agterberg et al. (2020). In the p-h channel, if $\bm{Q}=\bm{0}$, non-trivial QGN ($N^{\text{p-h},\bm{0}}\neq\mathbbm{1}$) leads to flavor-symmetry-breaking orders, including flavor ferromagnetism and inter-flavor coherent states; the $\bm{Q}\neq\bm{0}$ case promotes those orders to various forms of flavor DWs (e.g. spin DWs), with an additional possibility of pure charge DW order for $N^{\text{p-h},\bm{Q}}=\mathbbm{1}$.
@@ II Ideal interacting Hamiltonians
One caveat on the terminology should be noted: After projecting onto $\mathcal{H}_{\text{flat}}$, there are terms quadratic in fermion operators, besides the quartic ones that are conventionally called “interactions”. For the special case of Hubbard interactions, a property called the “uniform pairing condition” Tovmasyan et al. (2016) renders these terms trivial, and it was found that this condition can be guaranteed simply by certain point-group symmetries Herzog-Arbeitman et al. (2022). It is possible to generalize this condition, which we will exemplify in Sec. V.3. However, these quadratic terms are not of central focus since the current work is concerned with the existence of ideal interacting Hamiltonians for a given set of isolated flat bands, whose quantum geometry is encoded in the projection matrix and is unaffected by these terms to the leading order in $V/\Delta$. An operational workaround is to simply regard these additional quadratic terms as parts of $\hat{H}_{0}$. For more detailed discussions about this nomenclature subtlety, see Sec. II D of SM.
@@ III Solvable ground state
In the p-h case, the ground state of $\hat{H}_{\text{int}}$ can be similarly obtained by occupying any integer number of pseudo-bands of $\hat{\mathcal{E}}$ that are separated from other bands by finite gaps, since this gives rise to a non-degenerate  Note (3) and generically non-zero eigenvalue $\Phi_{0}$ within the corresponding particle number sector. This implies that the system is solvable only at certain rational filling fractions of the flat bands with specific values determined by the OP. It should be noted that when $\bm{Q}\neq 0$, $\hat{\mathcal{E}}$ has a folded BZ since it breaks the translation symmetry of the original electronic structure. Diagonalizing this pseudo-Hamiltonian then gives us $\mathsf{M}\mathsf{N}_{\text{flat}}$ pseudo-bands, where $\mathsf{M}$ is the smallest positive integer that makes $\mathsf{M}\bm{Q}=0$ modulo the reciprocal lattice. In contrast to the p-p case, here $\theta$ only cycles the pseudo-energies, permuting the equivalent translation-symmetry-breaking groundstates.
@@ IV.1 Few-body excitations
This is easier said than done for typical interacting Hamiltonians, where the commutator of a one-particle operator, $\gamma_{a}^{\dagger}$, with the Hamiltonian produces higher order terms such as $\gamma_{b}^{\dagger}\gamma_{c}^{\dagger}\gamma_{d}$, and the commutator of many-particle operators produces still higher order terms in a hierarchy that never closes. However, because $|\Psi\rangle$ is a zero-mode of all $\left(\hat{S}^{(I)}_{\bm{R}}-\langle\hat{S}^{(I)}_{\bm{R}}\rangle\right)$ within $\mathcal{H}_{\text{flat}}$, sectors with e.g. $\gamma_{a}^{\dagger}$ and $\gamma_{b}^{\dagger}\gamma_{c}^{\dagger}\gamma_{d}$ excitations cannot mix Kang and Vafek (2019); Bernevig et al. (2021); Herzog-Arbeitman et al. (2022); Vafek and Kang (2021) and thus the composition does not changed after the commutation. For example, for two-particle (Cooper pair) excitations,
@@
where $\bm{k}^{\pm}\equiv\bm{P}/2\pm\bm{k}$. The commutation relations for single-particle/hole, two-hole, and particle-hole (exciton) excitations can be similarly derived (see Sec. IV A-C of SM for computationally feasible expressions in terms of parameters in the Hamiltonian). By diagonalizing the scattering matrix $\Gamma$, one can compute the corresponding eigenstates and spectrum of an excitation Bernevig et al. (2021); Herzog-Arbeitman et al. (2022). For example, the charge $+1$ excitation spectrum only involves diagonalizing an effective single-particle Hamiltonian. These excitations can be computed with polynomial scaling in the system size, rather than exponentially as in most interacting systems, and in this sense are solvable. For certain cases when the form of interaction is simple (e.g. attractive Hubbard Herzog-Arbeitman et al. (2022)), the quasi-particle spectrum can be even analytically solvable.
@@ IV.2 Pseudo-spin $SU(2)$ symmetry of the p-p case and the phase stiffness
which quickly becomes comparable to $|\kappa|$ for non-zero symmetry-lowering energy scale $\delta$. We note that multiple factors can lead to a deviation from the idealized setup we constructed, and thus a weak explicit lowering of the pseudospin $SU(2)$ symmetry, including (but not limited to) the finite band gap, imperfect QGN and flatness of the bands, and the deviation of the interactions from the ideal form. Depending on the nature of those perturbations, the superconducting state may remain phase-coherent up to a temperature comparable to $|\kappa|$ for all dimensions $d\geq 2$. Our approach provides a direct estimate of the superfluid stiffness of nearly ideal systems, providing complementary estimates to generic upper bounds Hazra et al. (2019); Hofmann et al. (2022); Mao and Chowdhury (2023a, b).
@@ V.1 Spinful time-reversal symmetric flat bands
Then one can verify $N^{\text{p-p},\bm{Q}=\bm{0}}_{\mu s,\nu s^{\prime}}\equiv\sigma^{y}_{ss^{\prime}}\mathbbm{1}_{\mu\nu}$ satisfies the perfect QGN condition in the p-p channel with $\bm{Q}=\bm{0}$. Solvable Hamiltonians can thus be constructed with $B^{(I)}_{\mu s,\nu s^{\prime}}$ matrices in Eq. 15 taking the form $\sigma^{\mathsf{i}=0,x,y,z}_{ss^{\prime}}\tilde{B}^{(I)}_{\mu\nu}$, where $\sigma^{\mathsf{i}}$ are spin Pauli matrices and $\tilde{B}^{(I)}$ are arbitrary symmetric (for $\mathsf{i}=x,y,z$) or antisymmetric (for $\mathsf{i}=0$) hermitian matrices for the orbital indices. This includes a huge class of possible interactions, e.g. the attractive Hubbard interactions on each site and orbital, whose solvability was pointed out in Refs. Tovmasyan et al. (2016); Herzog-Arbeitman et al. (2022) and where the superfluid stiffness is shown to fundamentally depend on the flat-band quantum geometry and topology in several previous works Hofmann et al. (2023); Liang et al. (2017a); Verma et al. (2021); Hazra et al. (2018); Xie et al. (2020); Herzog-Arbeitman et al. (2022); Huhtinen et al. (2022); Herzog-Arbeitman et al. (2022).

### v3

@@ Abstract
The properties of a correlated electronic system are largely determined by its electronic structure. One such familiar manifestation is the role of Fermi surface nesting conditions Lomer (1962); Fedders and Martin (1966): if one portion of a Fermi surface overlaps with another under a shift by a wavevector $\bm{Q}$ (i.e. $\bm{k}\to\bm{k}+\bm{Q}$), the Fermi surface is said to be “perfectly nested” and consequently the system is prone to density wave (DW) order at $\bm{Q}$ with particle-hole (p-h) fermion bilinear order. If the overlap occurs under a momentum-space inversion about a certain wavevector $\bm{Q}/2$ (i.e. $\bm{k}\to\bm{Q}-\bm{k}$), the favored order is superconductivity (SC) with particle-particle (p-p) fermion bilinear order and a Cooper pair center of mass momentum $\bm{Q}$. (Such perfect nesting always exists at $\bm{Q}=0$ in time-reversal symmetric systems, where it leads to the familiar Cooper instability.) The thinking behind Fermi surface nesting is based on weak coupling mean-field analysis, in which ordering tendencies are diagnosed by divergent bare susceptibilities in corresponding channels. These concepts are indeed convenient and powerful (see e.g. Refs. Whangbo et al. (1991); Chuang et al. (2001); Kivelson et al. (2003); Shen et al. (2005); Lee et al. (2006); Johannes and Mazin (2008); Terashima et al. (2009); Ran et al. (2009); Nandkishore et al. (2012); Feng et al. (2013); Lee (2014); Jin et al. (2022) for typical applications). However, their validity is limited to weakly coupled systems in which the interaction strength is small compared to the bandwidth. In general, all forms of non-interacting susceptibilities grow as decreasing $T$ with the same asymptotic behavior $\chi\sim 1/T$ for $T$ greater than the bandwidth, which thus indicates comparable ordering tendencies for all possible order parameters. Significantly, this concept, defined solely by the electronic dispersion, overlooks another important piece of information in the electronic structure - the quantum geometry Provost and Vallee (1980); Resta (2011) and topology Bradlyn et al. (2017); Po et al. (2017); Kruthoff et al. (2017) encoded in the single-particle wavefunctions. An entirely different concept of “nesting” is thus needed for flat-band systems, whose bandwidths are small compared to the interaction strengths but which often have non-trivial quantum geometry and rich ordering phenomena Peotta and Törmä (2015); Tovmasyan et al. (2016); Kang and Vafek (2019); Rhim and Yang (2019); Hofmann et al. (2020); Bultinck et al. (2020); Kwan et al. (2021a); Khalaf et al. (2021); Lian et al. (2021); Bernevig et al. (2021); Ledwith et al. (2021); Kwan et al. (2021b); Vafek and Kang (2021); Peri et al. (2021); Liu et al. (2021); Rossi (2021); Angeli and MacDonald (2021); Wang et al. (2022); Hu et al. (2022); Huhtinen et al. (2022); Törmä et al. (2022); Chan et al. (2022a); Herzog-Arbeitman et al. (2022a); Mera and Mitscherling (2022); Pan et al. (2022); Zhang et al. (2022); Herzog-Arbeitman et al. (2022b); Lin et al. (2022); Herzog-Arbeitman et al. (2022a, 2024a); Peotta (2022); Khalaf et al. (2022); Chan et al. (2022b); Kang and Vafek (2019); Hofmann et al. (2023); Wan et al. (2023); Swaminathan et al. (2023); Hu et al. (2023); Löthman and Black-Schaffer (2017); Chen and Law (2024); Iskin (2023a); Peotta et al. (2023); Iskin (2023b); Tam and Peotta (2023); Zhang et al. (2023); Pan et al. (2023); Crépel et al. (2024); Mao et al. (2023); Kwan et al. (2024); Sahay et al. (2023); Jiang and Barlas (2023); Scheer and Lian (2023a, b); Devakul et al. (2023); Herzog-Arbeitman et al. (2024b); Jia et al. (2024); Wu et al. (2023); Chen and Huang (2023); Salerno et al. (2023); Huang et al. (2022); Yu et al. (2023); Kashihara et al. (2023); Soldini et al. (2023); Jiang et al. (2023); Yu et al. (2024); Peralta Gavensky et al. (2023); Setty et al. (2024); Verma et al. (2023); Setty et al. (2023a).
@@
Indeed, except for symmetry-enforced cases, it is unlikely for a physical system to have exact QGN - in common with the case of Fermi surface nesting. However, the use of QGN can be much more versatile: for systems with almost perfect QGN (quantification of the closeness to the ideal nesting will be defined below as ‘nestability’), one may engineer an idealized electronic structure with perfect QGN and then construct ideal interacting Hamiltonians that are close to the physical one. The solvability of these ideal models thus makes them excellent starting points for including perturbative corrections to reach the physical Hamiltonian. This type of intrinsically strong coupling viewpoint (in the sense that the dominant part of the interactions has been accounted for in the solvable models) has proven fruitful in recent studies of twisted bilayer graphene and other flat-band materials Tovmasyan et al. (2016); Kwan et al. (2021b); Vafek and Kang (2021); Kang and Vafek (2019); Khalaf et al. (2021); Bultinck et al. (2020); Lian et al. (2021); Bernevig et al. (2021); Ledwith et al. (2021); Pan et al. (2022); Lin et al. (2022); Herzog-Arbeitman et al. (2022a); Zhang et al. (2022); Herzog-Arbeitman et al. (2022b); Crépel et al. (2024); Zhang et al. (2023); Pan et al. (2023); Mao et al. (2023); Kwan et al. (2024). It is noteworthy that perfect QGN is an underlying feature of most of the known examples of solvable models for flat-band systems featuring fermion bi-linear order. We hence expect that QGN can greatly facilitate studies within this paradigm, in a way similar to the roles played by ‘ideal band conditions’ Roy (2014); Claassen et al. (2015); Mera and Ozawa (2021); Ledwith et al. (2022); Wang et al. (2021); Wang and Liu (2022); Wang et al. (2023); Dong et al. (2023) and “vortexability” Ledwith et al. (2023) in studies of fractional Chern insulator states.
@@ I.2 General Definitions
Depending on the channel, momentum $\bm{Q}$, and nesting matrix $N^{\bm{Q}}$, Eq. 8 or 9 represent fermion bilinear orders with different natures. In the p-p channel, if $\bm{Q}=\bm{0}$, the order is a uniform SC order; while for $\bm{Q}\neq\bm{0}$ it is pair-density-wave (PDW, not to be confused with the other DWs in the p-h channel) order of the single-$\bm{Q}$, Fulde–Ferrell type Agterberg et al. (2020). In the p-h channel, if $\bm{Q}=\bm{0}$, non-trivial QGN ($N^{\text{p-h},\bm{0}}\neq\mathbbm{1}$) leads to flavor-symmetry-breaking orders, including flavor ferromagnetism and inter-flavor coherent states; the $\bm{Q}\neq\bm{0}$ cases include various forms of DWs (e.g., spin, charge, or current/flux DWs).
@@ II Ideal interacting Hamiltonians
One caveat on the terminology should be noted: After projecting onto $\mathcal{H}_{\text{flat}}$, there are terms quadratic in fermion operators, besides the quartic ones that are conventionally called “interactions”. For the special case of Hubbard interactions, a property called the “uniform pairing condition” Tovmasyan et al. (2016) renders these terms trivial, and it was found that this condition can be guaranteed simply by certain point-group symmetries Herzog-Arbeitman et al. (2022a). It is possible to generalize this condition, which we will exemplify in Sec. V.3. However, these quadratic terms are not of central focus since the current work is concerned with the existence of ideal interacting Hamiltonians for a given set of isolated flat bands, whose quantum geometry is encoded in the projection matrix and is unaffected by these terms to the leading order in $V/\Delta$. An operational workaround is to simply regard these additional quadratic terms as parts of $\hat{H}_{0}$. For more detailed discussions about this nomenclature subtlety, see Sec. II D of SM.
@@ III Solvable ground state
In the p-h case, the ground state of $\hat{H}_{\text{int}}$ can be similarly obtained by occupying any integer number of pseudo-bands of $\hat{\mathcal{E}}$ that are separated from other bands by finite gaps, since this gives rise to a non-degenerate  Note (3) and generically non-zero eigenvalue $\Phi_{0}$ within the corresponding particle number sector. This implies that the system is solvable only at certain rational filling fractions of the flat bands with specific values determined by the OP. In contrast to the p-p case, here $\theta$ only cycles the pseudo-energies, permuting the equivalent translation-symmetry-breaking groundstates. It should be noted that when $\bm{Q}\neq 0$, $\hat{\mathcal{E}}$ has a folded BZ since it breaks the translation symmetry of the original electronic structure (corresponding to DW). Diagonalizing this pseudo-Hamiltonian then gives us $\mathsf{M}\mathsf{N}_{\text{flat}}$ pseudo-bands, where $\mathsf{M}$ is the smallest positive integer that makes $\mathsf{M}\bm{Q}=0$ modulo the reciprocal lattice. It is, in principle, possible to have $\varpi^{\bm{Q}}_{0}\approx 0$ in a continuous range of $\bm{Q}$, such that an incommensurate DW order is suggested. In such a case, the properties of each incommensurate $\bm{Q}$ should be approached to a good approximation from a nearby commensurate value.
@@
It should be noted that despite the spiritual similarity to the mean field theories, the Slater determinant states solved by the above approach are the exact many-body ground states of the ideal systems instead of just proxy states capturing the qualitative properties. The perturbations away from those ideal limits can be accounted for systematically since not only the ground states but also many aspects of the excitation spectrum can be solved, as we discuss below.
@@ IV.1 Few-body excitations
This is easier said than done for typical interacting Hamiltonians, where the commutator of a one-particle operator, $\gamma_{a}^{\dagger}$, with the Hamiltonian produces higher order terms such as $\gamma_{b}^{\dagger}\gamma_{c}^{\dagger}\gamma_{d}$, and the commutator of many-particle operators produces still higher order terms in a hierarchy that never closes. However, because $|\Psi\rangle$ is a zero-mode of all $\left(\hat{S}^{(I)}_{\bm{R}}-\langle\hat{S}^{(I)}_{\bm{R}}\rangle\right)$ within $\mathcal{H}_{\text{flat}}$, sectors with e.g. $\gamma_{a}^{\dagger}$ and $\gamma_{b}^{\dagger}\gamma_{c}^{\dagger}\gamma_{d}$ excitations cannot mix Kang and Vafek (2019); Bernevig et al. (2021); Herzog-Arbeitman et al. (2022a); Vafek and Kang (2021) and thus the composition does not changed after the commutation. For example, for two-particle (Cooper pair) excitations,
@@
where $\bm{k}^{\pm}\equiv\bm{P}/2\pm\bm{k}$. The commutation relations for single-particle/hole, two-hole, and particle-hole (exciton) excitations can be similarly derived (see Sec. IV A-C of SM for computationally feasible expressions in terms of parameters in the Hamiltonian). By diagonalizing the scattering matrix $\Gamma$, one can compute the corresponding eigenstates and spectrum of an excitation Bernevig et al. (2021); Herzog-Arbeitman et al. (2022a). For example, the charge $+1$ excitation spectrum only involves diagonalizing an effective single-particle Hamiltonian. These excitations can be computed with polynomial scaling in the system size, rather than exponentially as in most interacting systems, and in this sense are solvable. For certain cases when the form of interaction is simple (e.g. attractive Hubbard Herzog-Arbeitman et al. (2022a)), the quasi-particle spectrum can be even analytically solvable.
@@ IV.2 Pseudo-spin $SU(2)$ symmetry of the p-p case and the phase stiffness
which quickly becomes comparable to $|\kappa|$ for non-zero symmetry-lowering energy scale $\delta$. We note that multiple factors can lead to a deviation from the idealized setup we constructed, and thus a weak explicit lowering of the pseudospin $SU(2)$ symmetry, including (but not limited to) the finite band gap, imperfect QGN and flatness of the bands, and the deviation of the interactions from the ideal form. Depending on the nature of those perturbations, the superconducting state may remain phase-coherent up to a temperature comparable to $|\kappa|$ for all dimensions $d\geq 2$. Our approach provides a direct estimate of the superfluid stiffness of nearly ideal systems, providing complementary estimates to generic upper bounds Hazra et al. (2019a); Hofmann et al. (2022); Mao and Chowdhury (2023, 2024).
@@ V Examples of electronic structures with perfect QGN
Here, we present several examples of QGN. More detailed discussions, as well as another constructed model for a current density wave, can be found in Sec. V of the SM.
@@ V.1 Spinful time-reversal symmetric flat bands
Then one can verify $N^{\text{p-p},\bm{Q}=\bm{0}}_{\mu s,\nu s^{\prime}}\equiv\sigma^{y}_{ss^{\prime}}\mathbbm{1}_{\mu\nu}$ satisfies the perfect QGN condition in the p-p channel with $\bm{Q}=\bm{0}$. Solvable Hamiltonians can thus be constructed with $B^{(I)}_{\mu s,\nu s^{\prime}}$ matrices in Eq. 15 taking the form $\sigma^{\mathsf{i}=0,x,y,z}_{ss^{\prime}}\tilde{B}^{(I)}_{\mu\nu}$, where $\sigma^{\mathsf{i}}$ are spin Pauli matrices and $\tilde{B}^{(I)}$ are arbitrary symmetric (for $\mathsf{i}=x,y,z$) or antisymmetric (for $\mathsf{i}=0$) hermitian matrices for the orbital indices. This includes a huge class of possible interactions, e.g. the attractive Hubbard interactions on each site and orbital, whose solvability was pointed out in Refs. Tovmasyan et al. (2016); Herzog-Arbeitman et al. (2022a) and where the superfluid stiffness is shown to fundamentally depend on the flat-band quantum geometry and topology in several previous works Hofmann et al. (2023); Liang et al. (2017a); Verma et al. (2021); Hazra et al. (2019b); Xie et al. (2020); Herzog-Arbeitman et al. (2022b); Huhtinen et al. (2022); Herzog-Arbeitman et al. (2022a).